# Schedule On the Fly: Diffusion Time Prediction for Faster and Better Image Generation

### v2

@@ 4.1 Implementation details
We collect prompts to train our model. These prompts are generated with Florence-2 [61] and Llava-Next [28] to generate captions for the Laion-Art [54] and COYO-700M [5] datasets, and utilize these prompts to constitute our training set. We will elaborate on this in Appendix.
@@
We use the AdamW optimizer with $\beta_{1}$ = 0.9, $\beta_{2}$ = 0.99, a constant learning rate of 1e-5, and a maximum gradient norm of 1.0. Our TPM module typically requires only 200 sampled trajectories, 4 PPO epochs, and a batch size of 256 to effectively update its parameters.
@@ 4.2 Main Results
Dynamic Schedule for Different Images
In Fig. 4, we present images generated with different prompts and their corresponding schedules predicted by TPDM. When prompting TPDM with shorter and simpler prompts, there exists fewer objects and details in the generated image, thus the diffusion time decreases faster and reaches 00 in relatively fewer steps. On the contrary, when longer and more complex prompts are provided, the model needs to generate more visual details. Therefore, the diffusion time decreases slower to generate delicate details. In this case TPDM requires more denoising steps in the generation process.
@@
Table 1: Evaluation of models across various benchmarks, best, second best
@@
$\gamma$ in Eq. 9 controls how the image reward discounts with more generation steps, influencing how fast diffusion time decays during the denoising process, thereby affecting the average number of denoising steps of our model.
As shown in Fig. 6, as we decrease $\gamma$ from 0.97 to 0.85, TPDM tends to decrease the diffusion time more rapidly, leading to fewer sampling steps, from 15.0 to 7.5. Additionally, when compared to the baseline (yellow line), TPDM (purple line) consistently achieves a significantly higher aesthetic score with the same number of inference steps, achieving a good balance between model efficiency and generation performance.
@@
Our method achieves superior ability in generating fine-grained details. Images generated by TPDM have a more realistic laptop keyboard than both images from SD3-Medium in Fig. 5C, and the results in Fig.5D also demonstrate that TPDM can generate images without errors. This may due to that TPDM can adjust the diffusion schedule according to the denoising quality.
@@ 4.3 Quantitative Results
We apply TPM on several state-of-the-art diffusion models, including Stable Diffusion 3 Medium, Stable Diffusion 3.5 Large, and Flux 1.0 dev, and demonstrate how TPM could enhance their performance. We primarily evaluate metrics from two kinds: The first is objective metrics including FID, alignment with the given prompts (CLIP-T), and human preference scores (Aesthetic Score v2, and HPSv2.1). The second is directly comparing images generated from different models via user study.
@@
We compared TPDM and the above-mentioned models in Tab. 1. While maintaining competitive performance. All these models can generate images in about half of their recommended steps on average.
Moreover, the metrics representing human preference improve the most. For example, by generating with just 15.28 steps on average, TPDM-SD3-Medium gets a 29.59 HPS score, +1.07 higher than Stable Diffusion 3 with similar steps, also +0.47 higher than the original 28-step results. This may attributes to the reward model we utilized to calculate advantage in optimization, leading to aesthetically pleasing images aligned with human preferences. Furthermore, since other scheduling-tuning methods only conduct experiments on classical diffusion models such as Stable Diffusion 1.5 [45], we conduct an apple-to-apple comparison with these algorithms as well, which is demonstrated in Appendix Table 4.
@@
To better reflect human attitudes towards these models, we conduct a user study by inviting volunteers to compare images generated from different models and select the one they prefer.
Specifically, for each prompt, we provide two images from SD3-Medium generated with 15 and 28 steps separately, and image generated from TPDM-SD3-Medium. We invited 15 volunteers to evaluate images generated from 50 prompts. The result is shown in Tab. 2, indicating that our model can generate images that better align human preferences.
@@
4.4 Ablation on Module Architectures
In this section, we ablate on the choice of the inputs of TPM. As shown in Tab. 3, taking features from both the first and last layer into TPM performs better than only taking either of them.
@@ 5 Conclusions and Limitations
In this paper, we introduce the Time Prediction Diffusion Model (TPDM), a text-to-image diffusion model that has flexible denoising schedulers for different individual prompts. By introducing a Time Prediction Module, we effectively train TPDM by reinforcement learning guided by a reward model. Based on the current leading diffusion model architecture proposed in Stable Diffusion 3 Medium, we train a strong MM-DiT-based TPDM that shows competitive quantitative performances on multiple text-to-image generation benchmarks.
Despite its promising performances, TPDM has some limitations. For instance, in this paper, we only design a relatively simple architecture for TPM. Whether and how to improve such a module to obtain better performances remains unexplored. Second, we froze the parameters of the original model and employed our training method to update the parameters of the diffusion model, leading to improved results. This approach warrants further exploration.
@@ A.1 Dataset Source
We curated prompts from image captions in two high-quality datasets, Laion-art [54]and COYO-11M [6]. The images in Laion-art are recaptioned with Florence-2 [61], while COYO-11M retained its original captions.
@@ A.2 Filtering Pipeline
Appendix B Prompts of Figure
B.1 Figure 1
@@
Method Inference Steps FID CLIP-T Aesthetic v2 Pick Score HPSv2.1
DPM-Solver++ [30] 10 21.08 0.314 5.219 21.30 24.00
DPM-Solver++ & AYS [46] 10 18.81 0.313 5.156 21.31 24.40
DPM-Solver++ & GITS [7] 10 18.32 0.313 5.108 21.19 23.72
DPM-Solver++ & TPM 9.890 18.31 0.313 5.118 21.32 24.29
Table 4: Experiments on Stable Diffusion v1.5 [45]. 9.890 is the average step TPM used.
Appendix C Applicaion on Denoising Diffusion Probabilistic Models (DDPMs)
Prior to the widespread use of flow models, diffusion models dominates image generation, such as DDPM [13] and LDM [45]. They demonstrate several difference from flow models during inference. First, they directly predict the noise ($\epsilon$) added to the data at each step, rather than predicting the velocity vector field. Second, both the diffusion time $t$ and the noise level $\sigma_{t}$ are typically discrete, unlike continuous time in flow models. Moreover, the noise level $\sigma_{t}$ does not exhibit a simple linear relationship with the diffusion time $t$.
In this section, we integrate TPM into Stable Diffusion 1.5 [45], demonstrating that TPM also works in DDPMs. When determining the next diffusion time, TPM predicts the Beta distribution as in flow models, and then quantizes it to the closest $t$ in all $T$ discrete steps, thereby obtaining the corresponding noise level $\sigma_{t}$.
We can generally express their relationship as follows:
$\beta_{t}=\left(\sqrt{\beta_{start}}+\frac{t}{T-1}\left(\sqrt{\beta_{end}}-\sqrt{\beta_{start}}\right)\right)^{2},$ (10)
$\alpha_{t}=1-\beta_{t},$ (11)
$\bar{\alpha}_{t}=\prod_{s=1}^{t}\alpha_{s},$ (12)
$\sigma_{t}=\sqrt{\frac{1-\bar{\alpha}_{t}}{\bar{\alpha}_{t}}},$ (13)
where $\beta_{t}$ is the variance of the noise added at diffusion time $t$. $\alpha_{t}$ is the proportion of the original signal remaining. $\sigma_{t}$ is the total noise level, and $T$ is the total number of diffusion steps while training, typically set to 1000. $\beta_{start}$ and $\beta_{end}$ are both hyper-parameter, controlling the initial and final noise variances.
We compare the performance of TPM with other scheduled fine-tuning methods applied to diffusion models. The evaluation metrics are calculated using 5,000 prompts from the COCO 2017 validation set, adhering to the evaluation protocol outlined in Table 1. Notably, our findings reveal that TPM can be effectively combined with higher-order solvers, such as DPM-Solver++ [30], resulting in a significant improvement of -2.79 in FID, surpassing the performance of other methods. It is important to note that GITS [7] requires access to the statistics of the evaluation set. While this approach is advantageous for large-scale image generation, it may offer limited benefits when generating a single image. Despite that GITS makes use of additional information, TPM consistently outperforms GITS across all evaluation metrics.

### v3

@@ 4.1 Implementation details
We collect text prompts to train our model. These prompts are either from COCO 2017[24] train set or by captioning images in Laion-Art [54] and COYO-700M [5] datasets with Florence-2 [62] and Llava-Next [26]. We will elaborate on this in Appendix A.
@@
We utilize the AdamW optimizer with beta coefficients of $(0.9,0.99)$, a constant learning rate of $1\times 10^{-5}$, and a maximum gradient norm of 1.0. Our TPM module typically requires only 200 training steps. In each step, we sample a batch size of 256 trajectories, and update the model parameter 4 times with them.
Evaluation Metrics
We use FID, CLIP-T, Aesthetic v2, and Pick Score as evaluation metrics by testing with 5,000 prompts from COCO 2017 validation set. For HPS v2.1, we use the 3,200 prompts provided by the benchmark [60].
@@ 4.2 Main Results
Adaptive Schedule for Different Images
In Fig. 4, we present images generated with different prompts and their corresponding schedules predicted by TPDM. When prompting TPDM with shorter and simpler prompts, it requires fewer objects and details to appear in the generated image. Thus the diffusion time decreases faster and reaches 00 in relatively fewer steps. On the contrary, when more complex prompts are provided, the model needs to reason with more intermediate steps, so that it can generate more delicate details at early and middle stages. Therefore, the diffusion time decreases more slowly. In this case, TPDM requires more denoising steps in the generation process. In the Appendix C, we provide more examples of predicted schedules for analysis.
@@
Table 1: Evaluation of models across various benchmarks, The best result is highlighted in bold, and the second best result is underlined
@@
$\gamma$ in Eq. 9 controls how fast the reward decays when propagating from the final step to intermediate steps, as shown in Fig. 2, thereby affecting the average number of denosing steps of our converged TPDM.
As shown in Fig. 6, as we decrease $\gamma$ from 0.97 to 0.85, TPDM tends to decrease the diffusion time more rapidly, leading to fewer denosing steps, from 15.0 to 7.5. Additionally, when compared to the baseline (yellow line), TPDM (purple line) consistently achieves a significantly higher aesthetic score with the same number of inference steps.
@@
Fig.5 provides some specific images generated by both TPDM and SD3-Medium. In Fig. 5(c), the image generated by TPDM have a more realistic laptop keyboard than SD3-Medium with 15 and 28 steps. As for Fig.5(d), TPDM removes the extra baseball bat, leading to a more natural looking. Generally, TPDM is able to generate more accurate and realistic images.
@@ 4.3 Quantitative Results
We apply TPM on several state-of-the-art diffusion models, including SD3-Medium, SD3.5-Large, and Flux 1.0 dev, and demonstrate how TPM could enhance their performance. Other than commly used metrics like FID, CLIP-T and human preference scores (Aesthetic Score v2, and HPSv2.1), we provide a user study as well.
@@
We compared TPDM and the above-mentioned models in Tab. 1. While maintaining competitive performance, TPDM can generate images in about half of their recommended steps on average.
Among all the metrics, those representing human preference improve the most. For example, by generating with just 15.28 steps on average, TPDM-SD3-Medium gets a 29.59 HPS score, +1.07 higher than Stable Diffusion 3 with similar steps, also +0.47 higher than the original 28-step results. This improvement may be attributed to the reward model, which is well aligned with human preferences. By optimizing image generation based on this alignment, TPDM produces more aesthetically pleasing results. Furthermore, since other scheduler optimization methods mostly conduct experiments on classical diffusion models such as Stable Diffusion 1.5 [45], we conduct an apple-to-apple comparison with them in Appendix B as well.
@@
To better reflect human preference towards these models, we conduct a user study by inviting volunteers to compare images generated from different models.
Specifically, for each prompt, we provide two images generated by SD3-Medium with 15 and 28 steps separately, and an image generated by TPDM-SD3-Medium. We generate 50 groups of images for comparison in total, and invite 15 volunteers to select the ones they prefer. The result is shown in Tab. 2, indicating that our model can generate images that better align human preferences.
@@
Ablation on the Architecture
Tab. 3 indicates that taking features from both the first and last layer as the inputs for TPM performs better than only taking features from either the beginning or last two layers.
@@ 5 Conclusions and Limitations
In this paper, we view denoising steps as a kind of chain-of-thought in image generation, and introduce the Time Prediction Diffusion Model (TPDM) by adaptively predicting the next diffusion time to denoise images in the reverse diffusion process. This adjusts the noise schedule per sample, optimizing image generation for each prompt. By aligning the final outputs of the denoising process with human preferences, TPDM can reduce the number of inference steps by almost 50% while still keeping image quality.
Limitations. First, in this paper, we only design a simple architecture for TPM. A more delicate module may yield better performance. Second, we freeze the parameters of the original diffusion model. Updating them iteratively with the training of TPM could lead to improved results, which is left for further exploration.
Acknowledgement. This work was supported by National Natural Science Foundation of China under Grant No. 92467104, and Zhejiang Leading Innovative and Entrepreneur Team Introduction Program (2024R01007).
@@ A.1 Dataset Source
We curated training prompts from three high-quality datasets, COCO [24], Laion-art [54] and COYO-11M [6].
For COCO, the original captions in the training split is used. For Laion-art, we caption the images with Florence-2 [62] to obtain the text prompts. For COYO-11M, we retained its original captioning by Llava-next [26]. Fig. 7 illustrates the distribution of prompt length after tokenization.
Figure 7: Prompt length distribution.
@@ A.2 Filtering Pipeline
Method Inference Steps FID CLIP-T Aesthetic v2 Pick Score HPSv2.1
DPM-Solver++ [30] 10 21.08 0.314 5.219 21.30 24.00
DPM-Solver++ & AYS [46] 10 18.81 0.313 5.156 21.31 24.40
DPM-Solver++ & GITS [7] 10 18.32 0.313 5.108 21.19 23.72
DPM-Solver++ & TPM 9.89 18.31 0.313 5.118 21.32 24.29
Table 4: Experiments on Stable Diffusion v1.5 [45], and $9.89$ is the average number of steps TPM used. The best results are highlighted in bold.
Appendix B Applicaion on Denoising Diffusion Probabilistic Models (DDPMs)
Before the widespread use of flow-matching models, diffusion models dominated image generation, such as DDPM [13] and LDM [45]. They demonstrate several differences from flow models during inference. First, they directly predict the noise added to the data at each step, rather than predicting the velocity vector field. Second, both the diffusion time $t$ and the noise level are typically discrete, unlike continuous time in flow-matching models.
In this section, we integrate TPM into Stable Diffusion v1.5 [45], demonstrating that TPM also works with DDPMs. At diffusion time $t_{n-1}$, to determine the next diffusion time $t_{n}$, TPM predicts two real-valued parameters, $a_{n}$ and $b_{n}$. These parameters define $\alpha_{n}$ and $\beta_{n}$, which shape the Beta distribution of the diffusion time decay rate $r_{n}$, similar to the flow-matching model version of TPDM introduced in 3.2. Then $t_{n}$ is obtained using Eq. 6 and quantized to the nearest discrete diffusion time to obtain the corresponding noise level.
We compare the performance of TPDM with other scheduler optimization methods applied to diffusion models. The evaluation metrics are calculated using 5,000 prompts from the COCO 2017 validation set, in line with the evaluation protocol as in Table 1. Notably, our findings reveal that TPM can be combined with higher-order solvers, such as DPM-Solver++ [30], resulting in a significant improvement of -2.79 in FID, surpassing the performance of other methods as reported in Tab. 4.
Appendix C Analysis of the Predicted Schedule
Fig. 11 provides more examples to show the noise schedule predicted by TPDM.
We observe that TPDM tends to allocate more inference steps to higher noise levels to generate complex details and layouts. Take Fig. 11(a) for an example where multiple objects of various sizes are generated with a complex visual layout – TPDM allocates 10 out of 17 steps to more noisy diffusion time with $t>0.8$. This way, TPDM more efficiently spends its inference steps on denoising noisier samples at the earlier stage so that it can add more complex visual layouts and details as early as possible to eventually generate high quality results. It avoids the problem in the benchmark diffusion model that may waste many evenly allocated steps to denoise cleaner images at the later stage. The steep curve in Fig. 11(c) indicates that, for simple generation, TPDM reduces the diffusion time much faster to almost zero within only $13$ steps, instead of evenly allocating $28$ steps in the benchmark model. Fig. 11(e) and 11(f) also visualize the results for extremely long prompts. They exhibit a similar trend.
Moreover, Fig. 4 shows that simple prompts lead to faster decrease in diffusion time with fewer steps, while complex ones lead to more steps. We use ICNet [12] to score the image complexity, and calculate the average score of images generated with different number of steps in evaluation set. As shown in Fig. 10, it has a high correlation efficient (0.770), which supports our assumption that complex images are dynamically generated with more steps by TPDM.
Figure 10: Correlation between the number of steps and image complexity (measured by ICNet [12]).
Figure 11: Prompts, images and predicted schedules.
Appendix D Prompts to Generate Images in Fig. 1